use the following regression equation to estimate sales levels if price is set to $8: Sales = b0 + b1 Price What is the sales level when b0 = 100 and b1 = - 5
Answer #1

Sales = b0 + b1 Price

b0 = 100 and b1 = - 5

Sales = 100 -5*Price

Price = $8

Sales = 100 - 5*8

Sales = 100 - 40

Sales = 60